SEQUENTIAL CLUSTERING OF HCl IN ARGON: NEAR INFRARED SPECTRA OF Ar_{2}HCl AND Ar_{3}HCl

DAVID T. ANDERSON, \underline{SCOTT DAVIS} and DAVID J. NESBITT, JILA, National Institute of Standards and Technology and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0440.